A 1 mL sample of glycogen was calculated to contain 37 µmol (micromole) glucose. To 1 mL of this sample was added 2 mL of 2 M HCL. It was then hydrolysed by boiling the solution for 15 minutes. After boiling the hydrolysate was cooled and made up with H2O to a final volume of exactly 10 mL. The glucose was measured in this solution and found to have a concentration of 530 µg/mL (microgram/milliliter).

Assume the molecular mass of glucose is 180 g/mol and that the molecular mass of glucose within glycogen is 162 g/mol.

a) Calculate the mass (mg) of glucose in the 10 mL of hydrolysate. (As the 1 mL of glycogen sample was made up to a final volume of 10 mL, this mass of glucose was produced by the hydrolysis of the original 1 mL glycogen sample)??????????
1
Expert's answer
2017-05-31T09:58:11-0400
С = m/V
m = CV
M (C6H12O6) = 530 · 10 = 5300 µg = 5.3 mg